OpenStudy (anonymous):

Two negative integers with a sum of -12 and a product of 11. What are the integers?

OpenStudy (anonymous):

-11 and -1

OpenStudy (anonymous):

ok. thanks. another question. THe width of a rectangle is 2 inches less than the length. and the area is 35 square inches. What are the length and width of the rectangle? Need to find how to factor and then find solution set, finally length and width.

OpenStudy (anonymous):

\[A=wl=(l-2)(l)=35\]Now solve for the length:\[l^2-2l-35=0\]Use the quadatic formula:\[\frac{-(-2)^2\pm \sqrt{(-2)^2-4(1)(-35)}}{2(1)}\]This gives:\[\frac{-4\pm12}{2}\]We discard the negative solution as it is not physically relevant in this problem:\[l=4\]Now go back and solve for width:\[w=l-2=4-2=2\]This width is 2 inches and the length is 4 inches.

OpenStudy (anonymous):

oooops, messed up the quadratic formula lol. it is -b not -b^2 for the first term hehe. Answer should be (2(plus and minus)12)/2 so length is l=7 and width is 5 which makes sense since the are is 7*5=35 sorry for error